# Leveraging Online Olympiad-Level Math Problems for LLMs Training and Contamination-Resistant Evaluation

### v1

@@ Abstract
Advances in Large Language Models (LLMs) have sparked interest in their ability to solve Olympiad-level math problems. However, the training and evaluation of these models are constrained by the limited size and quality of available datasets, as creating large-scale data for such advanced problems requires extensive effort from human experts. In addition, current benchmarks are prone to contamination, leading to unreliable evaluations. In this paper, we present an automated pipeline that leverages the rich resources of the Art of Problem Solving (AoPS) forum, which predominantly features Olympiad-level problems and community-driven solutions. Using open-source LLMs, we develop a method to extract question-answer pairs from the forum, resulting in AoPS-Instruct, a dataset of more than 600,000 high-quality QA pairs. Our experiments demonstrate that fine-tuning LLMs on AoPS-Instruct improves their reasoning abilities across various benchmarks. Moreover, we build an automatic pipeline that introduces LiveAoPSBench, an evolving evaluation set with timestamps, derived from the latest forum data, providing a contamination-resistant benchmark for assessing LLM performance. Notably, we observe a significant decline in LLM performance over time, suggesting their success on older examples may stem from pre-training exposure rather than true reasoning ability. Our work presents a scalable approach to creating and maintaining large-scale, high-quality datasets for advanced math reasoning, offering valuable insights into the capabilities and limitations of LLMs in this domain. Our benchmark and code is available at https://github.com/DSL-Lab/aops.
@@ 1 Introduction
Using the most recent QA pairs, We build an automatic pipeline that introduces LiveAoPSBench, a contamination-resistant evaluation set for assessing the math reasoning capabilities of LLMs.
@@ 2 Related Work
Training Datasets for Math. Training datasets for mathematical reasoning can be categorized into two types: pretraining and supervised fine-tuning (SFT) datasets. First, pretraining datasets consist of large-scale math data, e.g., billions of tokens used during the pretraining phase of LLMs. Notable examples include Open-Web-Math (Paster et al., 2024) and Minerva (Lewkowycz et al., 2022), which contain 38.5B and 14.7B tokens of math data, respectively. Second, SFT datasets focus on high-quality question-answer pairs. Examples include Open-Math-Instruct (Toshniwal et al., 2024), Orca-math (Mitra et al., 2024), MetaMath (Yu et al., 2024), and the training sets of widely used benchmarks such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b). However, these datasets are generally limited to grade-school level mathematics and do not target more advanced topics like Olympiad-level math. One of the most closely related datasets to ours is Numina (Li et al., 2024), which combines popular SFT datasets like Orca-math, MATH, and GSM8K, along with approximately 190K new Olympiad-level QA pairs. Concurrently, Yue et al. (2024) introduced a large-scale instruction fine-tuning dataset for math and science, which has also shown improvements in mathematical reasoning. Table 1 presents a detailed comparison of our dataset with these related datasets.
@@
| Numina (Li et al., 2024) | K⋆ | K | ✗ | ✓ | GPT4-o |
| OpenMathInstruct (Toshniwal et al., 2024) | M | ✗ | ✗ | Mixtral† | |
| OlympiadBench He et al. (2024) | K | ✗ | ✓ | Human | |
| GSM8K (Cobbe et al., 2021) | K | K | ✗ | ✗ | Human |
| MATH (Hendrycks et al., 2021b) | K | K | ✗ | ✗ | Human |
| Orca-Math (Mitra et al., 2024) | K | - | ✗ | ✗ | GPT-4† |
@@
Contamination-Resistant Evaluation. Benchmarks that are publicly accessible are prone to be contaminated due to the potential inadvertent data overlap during training. The typical decontamination method involves using exact substring (e.g., -gram) matching to detect overlaps with the target evaluation sets (Zhuo et al., 2024). However, this approach fails to catch rephrased examples and can not eliminate all overlaps with the test set.(Yang et al., 2023). While alternative LLM-based methods for decontamination have been proposed, they often lack guarantees and may result in high false-positive rates (Yang et al., 2023). A reliable way to mitigate contamination is to select data that appeared after LLMs were trained, known as the knowledge cut-off. In the code generation domain, LiveCodeBench (Jain et al., 2024) addresses this issue by categorizing data based on timestamps, setting a cutoff date, and designating data beyond this point as unseen. We adopt a similar strategy in the math domain, partitioning the dataset by timestamps and enabling users to select data based on specific dates. Although this approach may not fully eliminate rephrased existing questions, it ensures that evaluation data remains unseen and less contaminated, providing a more accurate and fair assessment of LLMs.
@@ 4.3 Instruction Fine-Tuning
| Model | SFT Dataset | AoPS24 | Math | Olympiad Bench | Omni Math | AIME24 | AMC23 |
| No SFT | 11.7 | 47.1 | 14.5 | 12.3 | 1/30 | 8/40 | |
| Deepseek-Math | Numina | 16.3 | 55.5 | 22.7 | 17.0 | 0/30 | 12/40 |
| 7b-Ins | AoPS-Ins | 20.1 | 62.3 | 22.4 | 18.3 | 0/30 | 16/40 |
| Numina+AoPS-Ins | 19.7 | 58.8 | 25.6 | 18.0 | 2/30 | 11/40 | |
| No SFT | 15.40 | 56.30 | 21.20 | 15.90 | 0/30 | 16/40 | |
| Mathstral | Numina | 16.60 | 54.60 | 23.40 | 17.10 | 0/30 | 15/40 |
| 7B | AoPS-Ins | 23.60 | 60.80 | 27.10 | 19.90 | 0/30 | 12/40 |
| Numina+AoPS-Ins | 24.90 | 59.60 | 29.60 | 21.10 | 0/30 | 16/40 | |
| No SFT | 12.0 | 47.4 | 16.1 | 12.9 | 2/30 | 11/40 | |
| Llama-3.2 | Numina | 12.9 | 49.5 | 19.3 | 14.4 | 1/30 | 6/40 |
| 3B-Ins | AoPS-Ins | 17.1 | 52.9 | 18.5 | 15.1 | 0/30 | 11/40 |
| Numina+AoPS-Ins | 17.4 | 55.6 | 22.8 | 17.2 | 0/30 | 12/40 | |
| No SFT | 5.30 | 28.80 | 4.70 | 7.00 | 0/30 | 5/40 | |
| Llama-3.2 | Numina | 8.00 | 32.70 | 6.40 | 9.70 | 0/30 | 6/40 |
| 1B-Ins | AoPS-Ins | 10.00 | 34.70 | 11.10 | 11.00 | 0/30 | 6/40 |
| Numina+AoPS-Ins | 11.20 | 36.60 | 12.00 | 11.70 | 0/30 | 9/40 |
@@
Consistent with prior work, we train each model for three epochs (Shao et al., 2024; Yang et al., 2024b), as we observe additional epochs provide no further benefit (see Figure 9 in the Appendix for ablation studies on the number of epochs). We explore three data mixtures for fine-tuning: (1) AoPS alone, (2) Numina alone, and (3) AoPS + Numina. After fine-tuning each model, we evaluate the performance of each model on the following standard competition-level benchmarks: (1) OlympiadBench (He et al., 2024), which is an Olympiad-level evaluation dataset. Following prior literature (Yang et al., 2024a), we take only the math questions which have final answers and do not contain images or figures. This leaves us with samples from this dataset (2) Omni-MATH (Gao et al., 2024), which is a collection of problems from various mathematical olympiad competitions. (3) LiveAoPSBench set for the year 2024. The results are shown in Table 2. As shown by the table, fine-tuning with our dataset consistently boost the performance.
@@ 4.4 Ablation Studies
Evaluation Quality Assessment. We assess the quality of our evaluation set in two ways: by measuring its correlation with a well-established dataset and through manual evaluation over a subset of the data. First, He et al. (2024) compiled an Olympiad-level math evaluation set using manual assessment, which we leverage in our context to verify the quality of our method through the correlation between accuracies. Figure 5(a), demonstrates that the evaluation on LiveAoPSBench is highly correlated with carefully established benchmarks such as OlympiadBench. This demonstrates that our automatically generated benchmark aligns closely with the quality of those created through extensive human effort. Next, we subsample 10% of our evaluation set and ask human annotators to verify the correctness of the final parsed answers by referring to the original post. Annotators are given three options: yes, no, and no-answer. “Yes” and “no” indicate whether the answer is deemed correct, while “no-answer” is selected when a concrete answer is not appropriate (e.g., abstract concept questions answered with concrete examples). As a result, we found that 88% of the annotations were marked as correct, while 8% were incorrect and 4% fell under the no-answer category. To understand the gap from perfect accuracy here, we further measure the correlation between groups of human annotators by computing the percentage of choices that were consistent. Surprisingly, the human annotators only reached an agreement rate of 91%, demonstrating the challenge of evaluating Olympiad-level problems, even for graduate-level annotators.
@@ Appendix B Detailed Evaluation Results on LiveAoPSBench
B.1 Evaluating open-sourced LLMs
@@
B.2 Detailed Results
The accuracy comparison for these mainstream open source LLMs are shown in Tables 3, 4, 5 split by Month, Difficulty and Answer Type. The Month tables separately include evaluation results for 2023 and 2024. For the Difficulty and Answer Type tables, we use only the most recent evaluation results from 2024. Notably, the difficulty labels represent the general educational background of the problems rather than their exact difficulty. Over half of the problems originate from educational backgrounds associated with High School or High School Olympiads, and only around 7% are from Middle School, indicating our dataset’s focus is more on the complex problems. Similarly, in the Answer Type Table, more than half of the problems are categorized as numeric-int.
@@ C.1 Decontamination Details
We use 10-gram substring matching to decontaminate against test set for a comprehensive list of math evaluation datasets available. (Cobbe et al., 2021; Hendrycks et al., 2021b; He et al., 2024; AOPS, 2023; 2024; Zhang et al., 2023b; Lewkowycz et al., 2022; Gao et al., 2024; Miao et al., 2020; Hendrycks et al., 2021a; Koncel-Kedziorski et al., 2016; Patel et al., 2021; Zhang et al., 2023a). In Figure 8. We show the decontamination statistic for our dataset and Numina.
@@ D.2 Rewriting model ablation
Appendix E Use of AoPS as a Data Source
@@
Appendix F Prompts
Appendix G Dataset Examples
We provide further examples of our dataset and its rewritten solutions in Figures 13, 14, 15, 16. 17, and 18.

### v2

@@ Abstract
Advances in Large Language Models (LLMs) have sparked interest in their ability to solve Olympiad-level math problems. However, the training and evaluation of these models are constrained by the limited size and quality of available datasets, as creating large-scale data for such advanced problems requires extensive effort from human experts. In addition, current benchmarks are prone to contamination, leading to unreliable evaluations. In this paper, we present an automated pipeline that leverages the rich resources of the Art of Problem Solving (AoPS) forum, which predominantly features Olympiad-level problems and community-driven solutions. Using open-source LLMs, we develop a method to extract question-answer pairs from the forum, resulting in AoPS-Instruct, a dataset of more than 600,000 high-quality QA pairs. Our experiments demonstrate that fine-tuning LLMs on AoPS-Instruct improves their reasoning abilities across various benchmarks. Moreover, we build an automatic pipeline that introduces LiveAoPSBench, an evolving evaluation set with timestamps, derived from the latest forum data, providing a contamination-resistant benchmark for assessing LLM performance. Notably, we observe a significant decline in LLM performance over time, suggesting their success on older examples may stem from pre-training exposure rather than true reasoning ability. Our work presents a scalable approach to creating and maintaining large-scale, high-quality datasets for advanced math reasoning, offering valuable insights into the capabilities and limitations of LLMs in this domain. Our benchmark and code is available at https://livemathbench.github.io/leaderboard.html
@@ 1 Introduction
Using the most recent QA pairs, We build an automatic pipeline that introduces LiveAoPSBench, a contamination-resistant evaluation set for assessing the math reasoning capabilities of LLMs. LiveAoPSBench stays regularly updated.
@@ 2 Related Work
Training Datasets for Math. Training datasets for mathematical reasoning can be categorized into two types: pretraining and supervised fine-tuning (SFT) datasets. First, pretraining datasets consist of large-scale math data, e.g., billions of tokens used during the pretraining phase of LLMs. Notable examples include Open-Web-Math (Paster et al., 2024) and Minerva (Lewkowycz et al., 2022), which contain 38.5B and 14.7B tokens of math data, respectively. Second, SFT datasets focus on high-quality question-answer pairs. Examples include Open-Math-Instruct (Toshniwal et al., 2024), Orca-math (Mitra et al., 2024), MetaMath (Yu et al., 2024), DART-Math (Tong et al., 2024), ScaleQuest (Ding et al., 2024), (Shah et al., 2024), and the training sets of widely used benchmarks such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b). However, these datasets are generally limited to grade-school and intermediate high-school level mathematics and do not target more advanced topics like Olympiad-level math. One of the most closely related datasets to ours is Numina (Li et al., 2024), which combines popular SFT datasets like Orca-math, MATH, and GSM8K, along with approximately 190K new Olympiad-level QA pairs. Concurrently, Yue et al. (2024) introduced a large-scale instruction fine-tuning dataset for math and science, which has also shown improvements in mathematical reasoning. Table 1 presents a detailed comparison of our dataset with these related datasets.
@@
| Numina | K⋆ | K | ✗ | ✓ | GPT4-o |
| OpenMathInstruct | M | ✗ | ✗ | Mixtral† | |
| OlympiadBench | K | ✗ | ✓ | Human | |
| DART-Math | K | K | ✗ | ✗ | DeepSeekMath† |
| GSM8K | K | K | ✗ | ✗ | Human |
| MATH | K | K | ✗ | ✗ | Human |
| Orca-Math | K | - | ✗ | ✗ | GPT-4† |
| MAmmoTH2 | M | - | ✗ | ✗ | Mixtral |
@@
Contamination-Resistant Evaluation. Benchmarks that are publicly accessible are prone to be contaminated due to the potential inadvertent data overlap during training. The typical decontamination method involves using exact substring (e.g., -gram) matching to detect overlaps with the target evaluation sets (Zhuo et al., 2024). However, this approach fails to catch rephrased examples and can not eliminate all overlaps with the test set (Yang et al., 2023). While alternative LLM-based methods for decontamination have been proposed, they often lack guarantees and may result in high false-positive rates (Yang et al., 2023). A reliable way to mitigate contamination is to select data that appeared after LLMs were trained, known as the knowledge cut-off. In the code generation domain, LiveCodeBench (Jain et al., 2024) addresses this issue by categorizing data based on timestamps, setting a cutoff date, and designating data beyond this point as unseen. We adopt a similar strategy in the math domain, partitioning the dataset by timestamps and enabling users to select data based on specific dates. Although this approach may not fully eliminate rephrased existing questions, it ensures that evaluation data remains unseen and less contaminated, providing a more accurate and fair assessment of LLMs.
@@ 4.1 Dataset Statistics
4.2 Evaluation Set with Timestamps Mitigate Contamination
| Time | 23/01-04 | 23/05-08 | 23/09-12 | 24/01-04 | 24/05-08 |
| Overlap (%) | 13.24 | 11.65 | 12.82 | 9.92 | 6.88 |
| Count | 229 | 208 | 218 | 226 | 109 |
We analyze the relationship between timestamps and contamination using our evaluation set, tested against the Numnia Training Set (released July 2024). As shown in Table 2, we report contamination levels based on 10-gram overlap rates across different timestamps. We observe a consistent decrease in contamination rate as the test data timestamps become more recent.
@@ 4.4 Instruction Fine-Tuning
| Model | SFT Dataset | AoPS24 | Math | Olympiad Bench | Omni Math |
| No SFT | 11.7 | 47.1 | 14.5 | 12.3 | |
| Deepseek-Math | Numina | 16.3 | 55.5 | 22.7 | 17.0 |
| 7b-Ins | AoPS-Ins | 19.0 | 58.8 | 24.3 | 17.8 |
| Numina+AoPS-Ins | 19.7 | 58.8 | 25.6 | 18.0 | |
| No SFT | 15.40 | 56.30 | 21.20 | 15.90 | |
| Mathstral | Numina | 16.60 | 54.60 | 23.40 | 17.10 |
| 7B | AoPS-Ins | 23.60 | 60.80 | 27.10 | 19.90 |
| Numina+AoPS-Ins | 24.90 | 59.60 | 29.60 | 21.10 | |
| No SFT | 12.0 | 47.4 | 16.1 | 12.9 | |
| Llama-3.2 | Numina | 12.9 | 49.5 | 19.3 | 14.4 |
| 3B-Ins | AoPS-Ins | 16.7 | 54.6 | 19.6 | 16.4 |
| Numina+AoPS-Ins | 17.4 | 55.6 | 22.8 | 17.2 | |
| No SFT | 5.30 | 28.80 | 4.70 | 7.00 | |
| Llama-3.2 | Numina | 8.00 | 32.70 | 6.40 | 9.70 |
| 1B-Ins | AoPS-Ins | 10.00 | 34.70 | 11.10 | 11.00 |
| Numina+AoPS-Ins | 11.20 | 36.60 | 12.00 | 11.70 |
@@
Consistent with prior work, we train each model for three epochs (Shao et al., 2024; Yang et al., 2024b), as we observe additional epochs provide no further benefit (see Figure 10 in the Appendix for ablation studies on the number of epochs). We explore three data mixtures for fine-tuning: (1) AoPS alone, (2) Numina alone, and (3) AoPS + Numina. For each dataset, we use the full decontaminated dataset for fine-tuning (i.e., Numina with K QA pairs, and AoPS-Instruct with K QA pairs). After fine-tuning each model, we evaluate the performance of each model on the following standard competition-level benchmarks: (1) OlympiadBench (He et al., 2024), which is an Olympiad-level evaluation dataset. Following prior literature (Yang et al., 2024a), we take only the math questions which have final answers and do not contain images or figures. This leaves us with samples from this dataset (2) Omni-MATH (Gao et al., 2024), which is a collection of problems from various mathematical olympiad competitions. (3) LiveAoPSBench set for the year 2024. The results are shown in Table 3. As shown by the table, fine-tuning with our dataset consistently boost the performance.
@@ 4.5 Ablation Studies
Evaluation Quality Assessment. We assess the quality of our evaluation set in two ways: by measuring its correlation with a well-established dataset and through manual evaluation over a subset of the data. First, He et al. (2024) compiled an Olympiad-level math evaluation set using manual assessment, which we leverage in our context to verify the quality of our method through the correlation between accuracies. Figure 5(a), demonstrates that the evaluation on LiveAoPSBench is highly correlated with carefully established benchmarks such as OlympiadBench. This demonstrates that our automatically generated benchmark aligns closely with the quality of those created through extensive human effort. Next, we subsample 10% of our evaluation set and ask human annotators to verify the correctness of the final parsed answers by referring to the original post. Annotators are given three options: yes, no, and no-answer. “Yes” and “no” indicate whether the answer is deemed correct, while “no-answer” is selected when a concrete answer is not appropriate (e.g., abstract concept questions answered with concrete examples). As a result, we found that 92% of the annotations were marked as correct, while 5% were incorrect and 3% fell under the no-answer category.
@@ 6 Conclusion and Future Work
Impact Statement
This paper aims to advance the field of mathematical reasoning. It is important to note that AoPS-Instruct and LiveAoPSBench contain no personal information. Additionally, we have not identified any societal consequences that require special attention in this context.
Acknowledgments
This work was funded, in part, by the NSERC DG Grant (No. RGPIN-2022-04636), the Vector Institute for AI, Canada CIFAR AI Chair, NSERC Canada Research Chair (CRC), NSERC Discovery Grants, a Google Gift Fund, and the Government of Canada’s New Frontiers in Research Fund NFRFE-2023-00936. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through the Digital Research Alliance of Canada alliance.can.ca, and companies sponsoring the Vector Institute www.vectorinstitute.ai/#partners, and Advanced Research Computing at the University of British Columbia. Additional hardware support was provided by John R. Evans Leaders Fund CFI grant. Sadegh Mahdavi and Muchen Li are supported by UBC Four Year Doctoral Fellowships.
@@ Appendix B Detailed Evaluation Results on LiveAoPSBench
B.1 LiveAoPSBench-2024
We update the LiveAoPSBench with the complete 2024 data. This resulted in LiveAoPSBench-2024 with 5328 questions. In figure 8 we provide the latest result with the state-of-the-art reasoning models like deepseek-r1(DeepSeek-AI, 2025) and QwQ(Team, 2024).
B.2 Evaluating open-sourced LLMs
@@
B.3 Detailed Results
The accuracy comparison for these mainstream open source LLMs are shown in Tables 4, 5, 6 split by Month, Difficulty and Answer Type. The Month tables separately include evaluation results for 2023 and 2024. For the Difficulty and Answer Type tables, we use only the most recent evaluation results from 2024. Notably, the difficulty labels represent the general educational background of the problems rather than their exact difficulty. Over half of the problems originate from educational backgrounds associated with High School or High School Olympiads, and only around 7% are from Middle School, indicating our dataset’s focus is more on the complex problems. Similarly, in the Answer Type Table, more than half of the problems are categorized as numeric-int.
@@ C.1 Decontamination Details
We use 10-gram substring matching to decontaminate against test set for a comprehensive list of math evaluation datasets available. (Cobbe et al., 2021; Hendrycks et al., 2021b; He et al., 2024; AOPS, 2023, 2024; Zhang et al., 2023b; Lewkowycz et al., 2022; Gao et al., 2024; Miao et al., 2020; Hendrycks et al., 2021a; Koncel-Kedziorski et al., 2016; Patel et al., 2021; Zhang et al., 2023a). In Figure 9. We show the decontamination statistic for our dataset and Numina.
@@ D.2 Rewriting model ablation
Appendix E AIME and AMC results on the main models
Results of fine-tuning on AIME 2024 and AMC 2024 are present in table 8.
| Model | SFT Dataset | AIME24 | AMC23 |
| No SFT | 1/30 | 8/40 | |
| Deepseek-Math | Numina | 0/30 | 12/40 |
| 7b-Ins | AoPS-Ins | 1/30 | 15/40 |
| Numina+AoPS-Ins | 2/30 | 11/40 | |
| No SFT | 0/30 | 16/40 | |
| Mathstral | Numina | 0/30 | 15/40 |
| 7B | AoPS-Ins | 0/30 | 12/40 |
| Numina+AoPS-Ins | 0/30 | 16/40 | |
| No SFT | 2/30 | 11/40 | |
| Llama-3.2 | Numina | 1/30 | 6/40 |
| 3B-Ins | AoPS-Ins | 2/30 | 11/40 |
| Numina+AoPS-Ins | 0/30 | 12/40 | |
| No SFT | 0/30 | 5/40 | |
| Llama-3.2 | Numina | 0/30 | 6/40 |
| 1B-Ins | AoPS-Ins | 0/30 | 6/40 |
| Numina+AoPS-Ins | 0/30 | 9/40 |
| Dataset | Size | LiveAoPSBench | Olympiad Bench | Omni Math | MATH |
| DART-MATH-Hard | K | 14.4 | 21.8 | 15.4 | 52.5 |
| AoPS-Instruct | K | 19.0 | 24.3 | 17.8 | 58.8 |
Appendix F Use of AoPS as a Data Source
@@
Appendix G Prompts
Appendix H Dataset Examples
We provide further examples of our dataset and its rewritten solutions in Figures 14, 15, 16, 17. 18, and 19.